# Unifying Large Language Models and Knowledge Graphs: A Roadmap

### v1

@@ Abstract
Large language models (LLMs), such as ChatGPT and GPT4, are making new waves in the field of natural language processing and artificial intelligence, due to their emergent ability and generalizability. However, LLMs are black-box models, which often fall short of capturing and accessing factual knowledge. In contrast, Knowledge Graphs (KGs), Wikipedia and Huapu for example, are structured knowledge models that explicitly store rich factual knowledge. KGs can enhance LLMs by providing external knowledge for inference and interpretability. Meanwhile, KGs are difficult to construct and evolving by nature, which challenges the existing methods in KGs to generate new facts and represent unseen knowledge. Therefore, it is complementary to unify LLMs and KGs together and simultaneously leverage their advantages. In this article, we present a forward-looking roadmap for the unification of LLMs and KGs. Our roadmap consists of three general frameworks, namely, 1) KG-enhanced LLMs, which incorporate KGs during the pre-training and inference phases of LLMs, or for the purpose of enhancing understanding of the knowledge learned by LLMs; 2) LLM-augmented KGs, that leverage LLMs for different KG tasks such as embedding, completion, construction, graph-to-text generation, and question answering; and 3) Synergized LLMs + KGs, in which LLMs and KGs play equal roles and work in a mutually beneficial way to enhance both LLMs and KGs for bidirectional reasoning driven by both data and knowledge. We review and summarize existing efforts within these three frameworks in our roadmap and pinpoint their future research directions.
@@
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ 1 Introduction
Figure 1: Summarization of the pros and cons for LLMs and KGs. LLM pros: General Knowledge [11], Language Processing [12], Generalizability [13]; LLM cons: Implicit Knowledge [14], Hallucination [15], Indecisiveness [16], Black-box [17], Lacking Domain-specific/New Knowledge [18]. KG pros: Structural Knowledge [19], Accuracy [20], Decisiveness [21], Interpretability [22], Domain-specific Knowledge [23], Evolving Knowledge [24]; KG cons: Incompleteness [25], Lacking Language Understanding [26], Unseen Facts [27].
@@ 2.1 Large Language models (LLMs)
Large language modes (LLMs) pre-trained on large-scale corpus have shown great potential in various NLP tasks [13]. As shown in Fig. 3, most LLMs derive from the Transformer design [50], which contains the encoder and decoder modules empowered by a self-attention mechanism. Based on the architecture structure, LLMs can be categorized into three groups: 1) encoder-only LLMs2) encoder-decoder LLMs, and 3) decoder-only LLMs. As shown in Fig. 2, we summarize several representative LLMs with different model architectures, model sizes, and open-source availabilities.
@@ 2.2 Prompt Engineering
Prompt engineering seeks to improve the capacity of large large language models (e.g.,ChatGPT) in diverse complex tasks such as question answering, sentiment classification, and common sense reasoning. Chain-of-thought (CoT) prompt [64] enables complex reasoning capabilities through intermediate reasoning steps. Liu et al. [65] incorporate external knowledge to design better knowledge-enhanced prompts. Automatic prompt engineer (APE) proposes an automatic prompt generation method to improve the performance of LLMs [66]. Prompt offers a simple way to utilize the potential of LLMs without finetuning. Proficiency in prompt engineering leads to a better understanding of the strengths and weaknesses of LLMs.
@@ 2.3.3 Domain-specific Knowledge Graphs
AliOpenKG Recommendation https://shorturl.at/bmpDH
@@ 2.4 Applications
LLMs as KGs have been widely applied in various real-world applications. We summarize some representative applications of using LLMs and KGs in Table I. ChatGPT/GPT-4 are LLM-based chatbots that can communicate with humans in a natural dialogue format. To improve knowledge awareness of LLMs, ERNIE 3.0 and Bard incorporate KGs into their chatbot applications. Instead of Chatbot. Firefly develops a photo editing application that allows users to edit photos by using natural language descriptions. Copilot, New Bing, and Shop.ai adopt LLMs to empower their applications in the areas of coding assistant, web search, and recommendation, respectively. Wikidata and KO are two representative knowledge graph applications that are used to provide external knowledge. AliOpenKG is a knowledge graph designed for recommendation. Doctor.ai develops a health care assistant that incorporates LLMs and KGs to provide medical advice.
@@ 3.1 Roadmap
The roadmap of unifying KGs and LLMs is illustrated in Fig. 6. In the roadmap, we identify three frameworks for the unification of LLMs and KGs, including KG-enhanced LLMs, LLM-augmented KGs, and Synergized LLMs + KGs.
@@ 4 KG-enhanced LLMs
Task Method Year Technique
KG-enhanced LLM pre-training ERNIE [35] 2019 Integrating KGs into Training Objective
Skep [104] 2020 Integrating KGs into Training Objective
GLM [105] 2020 Integrating KGs into Training Objective
Ebert [106] 2020 Integrating KGs into Training Objective
KEPLER [40] 2021 Integrating KGs into Training Objective
Deterministic LLM [107] 2022 Integrating KGs into Training Objective
KALA [108] 2022 Integrating KGs into Training Objective
WKLM [109] 2020 Integrating KGs into Training Objective
K-BERT [36] 2020 Integrating KGs into Language Model Inputs
CoLAKE [110] 2020 Integrating KGs into Language Model Inputs
ERNIE3.0 [103] 2021 Integrating KGs into Language Model Inputs
DkLLM [111] 2022 Integrating KGs into Language Model Inputs
Dict-BERT [112] 2022 Integrating KGs into Language Model Inputs
BERT-MK [113] 2020 Integrating KGs into Additional Fusion Modules
Jaket [114] 2020 Integrating KGs into Additional Fusion Modules
K-Adapter [115] 2020 Integrating KGs into Additional Fusion Modules
cokebert [116] 2021 Integrating KGs into Additional Fusion Modules
KG-enhanced LLM inference QA-GNN [117] 2021 Dynamic Knowledge Fusion
Jointlk [118] 2022 Dynamic Knowledge Fusion
GreaseLM [119] 2022 Dynamic Knowledge Fusion
KGLM [120] 2019 Retrival-augmented knowledge fusion
REALM [121] 2020 Retrival-augmented knowledge fusion
RAG [93] 2020 Retrival-augmented knowledge fusion
Story-fragments [122] 2021 Retrival-augmented knowledge fusion
EMAT [123] 2022 Retrival-augmented knowledge fusion
KG-enhanced LLM interpretability LAMA [14] 2019 KGs for LLM probing
LPAQA [124] 2020 KGs for LLM probing
Autoprompt [125] 2020 KGs for LLM probing
Adolphs et al. [126] 2021 KGs for LLM probing
MedLAMA [127] 2022 KGs for LLM probing
Alex et al. [128] 2022 KGs for LLM probing
KagNet [38] 2019 KGs for LLM analysis
interpret-lm [129] 2021 KGs for LLM analysis
knowledge-neurons [39] 2021 KGs for LLM analysis
Shaobo et al. [130] 2022 KGs for LLM analysis
@@ 4.1 KG-enhanced LLM Pre-training
Existing large language models mostly rely on unsupervised training on the large-scale corpus. While these models may exhibit impressive performance on downstream tasks, they often lack practical knowledge relevant to the real world. Previous works that integrate KGs into large language models can be categorized into three parts: 1) Integrating KGs into training objective, 2) Integrating KGs into LLM inputs and 3) Integrating KGs into additional fusion modules.
@@ 4.1.1 Integrating KGs into Training Objective
The other line of work explicitly leverages the connections with knowledge and input text. As shown in Figure 9, ERNIE [91] proposes a novel word-entity alignment training objective as a pre-training objective. Specifically, ERNIE feeds both sentences and corresponding entities mentioned in the text into LLMs, and then trains the LLMs to predict alignment links between textual tokens and entities in knowledge graphs. Similarly, KALM [92] enhances the input tokens by incorporating entity embeddings and includes an entity prediction pre-training task in addition to the token-only pre-training objective. This approach aims to improve the ability of LLMs to capture knowledge related to entities. Finally, KEPLER [131] directly employs both knowledge graph embedding training objective and Masked token pre-training objective into a shared transformer-based encoder. Deterministic LLM [107] focuses on pre-training language models to capture deterministic factual knowledge. It only masks the span that has a deterministic entity as the question and introduces additional clue contrast learning and clue classification objective. WKLM [109] first replaces entities in the text with other same-type entities and then feeds them into LLMs. The model is further pre-trained to distinguish whether the entities have been replaced or not.
@@ 4.1.2 Integrating KGs into LLM Inputs
The above methods can indeed inject a large amount of knowledge into LLMs. However, they mostly focus on popular entities and overlook the low-frequent and long-tail ones. DkLLM [111] aims to improve the LLMs representations towards those entities. DkLLM first proposes a novel measurement to determine long-tail entities and then replaces these selected entities in the text with pseudo token embedding as new input to the large language models. Furthermore, Dict-BERT [112] proposes to leverage external dictionaries to solve this issue. Specifically, Dict-BERT improves the representation quality of rare words by appending their definitions from the dictionary at the end of input text and trains the language model to locally align rare words representations in input sentences and dictionary definitions as well as to discriminate whether the input text and definition are correctly mapped.
@@
4.1.3 Integrating KGs by Additional Fusion Modules
By introducing additional fusion modules into LLMs, the information from KGs can be separately processed and fused into LLMs. As shown in Fig. 11, ERNIE [91] proposes a textual-knowledge dual encoder architecture where a T-encoder first encodes the input sentences, then a K-encoder processes knowledge graphs which are fused them with the textual representation from the T-encoder. BERT-MK [113] employs a similar dual-encoder architecture but it introduces additional information of neighboring entities in the knowledge encoder component during the pre-training of LLMs. However, some of the neighboring entities in KGs may not be relevant to the input text, resulting in extra redundancy and noise. CokeBERT [116] focuses on this issue and proposes a GNN-based module to filter out irrelevant KG entities using the input text. JAKET [114] proposes to fuse the entity information in the middle of the large language model. The first half of the model processes the input text and knowledge entity sequence separately. Then, the outputs of text and entities are combined together. Specifically, the entity representations are added to their corresponding position of text representations, which are further processed by the second half of the model. K-adapters [115] fuses linguistic and factual knowledge via adapters which only adds trainable multi-layer perception in the middle of the transformer layer while the existing parameters of large language models remain frozen during the knowledge pre-training stage. Such adapters are independent of each other and can be trained in parallel.
Figure 11: Integrating KGs into LLMs by additional fusion modules.
@@ 4.2 KG-enhanced LLM Inference
The above methods could effectively fuse knowledge with the textual representations in the large language models. However, real-world knowledge is subject to change and the limitation of these approaches is that they do not permit updates to the incorporated knowledge without retraining the model. As a result, they may not generalize well to the unseen knowledge during inference [132]. Therefore, considerable research has been devoted to keeping the knowledge space and text space separate and injecting the knowledge while inference. These methods mostly focus on the Question Answering (QA) tasks, because QA requires the model to capture both textual semantic meanings and up-to-date real-world knowledge.
4.2.1 Dynamic Knowledge Fusion
A straightforward method is to leverage a two-tower architecture where one separated module processes the text inputs and the other one processes the relevant knowledge graph inputs [133]. However, this method lacks interaction between text and knowledge. Thus, KagNet [94] proposes to first encode the input KG, and then augment the input textual representation. In contrast, MHGRN [134] uses the final LLM outputs of the input text to guide the reasoning process on the KGs. Yet, both of them only design a single-direction interaction between the text and KGs. To tackle this issue, QA-GNN [117] proposes to use a GNN-based model to jointly reason over input context and KG information via message passing. Specifically, QA-GNN represents the input textual information as a special node via a pooling operation and connects this node with other entities in KG. However, the textual inputs are only pooled into a single dense vector, limiting the information fusion performance. JointLK [118] then proposes a framework with fine-grained interaction between any tokens in the textual inputs and any KG entities through LM-to-KG and KG-to-LM bi-directional attention mechanism. As shown in Fig. 12, pairwise dot-product scores are calculated over all textual tokens and KG entities, the bi-directional attentive scores are computed separately. In addition, at each jointLK layer, the KGs are also dynamically pruned based on the attention score to allow later layers to focus on more important sub-KG structures. Despite being effective, in JointLK, the fusion process between the input text and KG still uses the final LLM outputs as the input text representations. GreaseLM [119] designs deep and rich interaction between the input text tokens and KG entities at each layer of the LLMs. The architecture and fusion approach are mostly similar to ERNIE [91] discussed in Section 4.1.3, except that GreaseLM does not use the text-only T-encoder to handle input text.
Figure 12: Dynamic knowledge graph fusion for LLM inference.
@@ 4.2.2 Retrieval-Augmented Knowledge Fusion
Different from the above methods that store all knowledge in parameters, as shown in Figure 13, RAG [93] proposes to combine non-parametric and parametric modules to handle the external knowledge. Given the input text, RAG first searches for relevant KG in the non-parametric module via MIPS to obtain several documents. RAG then treats these documents as hidden variables $z$ and feeds them into the output generator, empowered by Seq2Seq LLMs, as additional context information. The research indicates that using different retrieved documents as conditions at different generation steps performs better than only using a single document to guide the whole generation process. The experimental results show that RAG outperforms other parametric-only and non-parametric-only baseline models in open-domain QA. RAG can also generate more specific, diverse, and factual text than other parameter-only baselines. Story-fragments [122] further improves architecture by adding an additional module to determine salient knowledge entities and fuse them into the generator to improve the quality of generated long stories. EMAT [123] further improves the efficiency of such a system by encoding external knowledge into a key-value memory and exploiting the fast maximum inner product search for memory querying. REALM [121] proposes a novel knowledge retriever to help the model to retrieve and attend over documents from a large corpus during the pre-training stage and successfully improves the performance of open-domain question answering. KGLM [120] selects the facts from a knowledge graph using the current context to generate factual sentences. With the help of an external knowledge graph, KGLM could describe facts using out-of-domain words or phrases.

### v3

@@ Abstract
Large language models (LLMs), such as ChatGPT and GPT4, are making new waves in the field of natural language processing and artificial intelligence, due to their emergent ability and generalizability. However, LLMs are black-box models, which often fall short of capturing and accessing factual knowledge. In contrast, Knowledge Graphs (KGs), Wikipedia and Huapu for example, are structured knowledge models that explicitly store rich factual knowledge. KGs can enhance LLMs by providing external knowledge for inference and interpretability. Meanwhile, KGs are difficult to construct and evolve by nature, which challenges the existing methods in KGs to generate new facts and represent unseen knowledge. Therefore, it is complementary to unify LLMs and KGs together and simultaneously leverage their advantages. In this article, we present a forward-looking roadmap for the unification of LLMs and KGs. Our roadmap consists of three general frameworks, namely, 1) KG-enhanced LLMs, which incorporate KGs during the pre-training and inference phases of LLMs, or for the purpose of enhancing understanding of the knowledge learned by LLMs; 2) LLM-augmented KGs, that leverage LLMs for different KG tasks such as embedding, completion, construction, graph-to-text generation, and question answering; and 3) Synergized LLMs + KGs, in which LLMs and KGs play equal roles and work in a mutually beneficial way to enhance both LLMs and KGs for bidirectional reasoning driven by both data and knowledge. We review and summarize existing efforts within these three frameworks in our roadmap and pinpoint their future research directions.
@@
publicationid: pubid: $0$–$00$ 2023 IEEE
@@ 1 Introduction
Figure 1: Summarization of the pros and cons for LLMs and KGs. LLM pros: General Knowledge [11], Language Processing [12], Generalizability [13]; LLM cons: Implicit Knowledge [14], Hallucination [15], Indecisiveness [16], Black-box [17], Lacking Domain-specific/New Knowledge [18]. KG pros: Structural Knowledge [19], Accuracy [20], Decisiveness [21], Interpretability [22], Domain-specific Knowledge [23], Evolving Knowledge [24]; KG cons: Incompleteness [25], Lacking Language Understanding [26], Unseen Facts [27]. Pros. and Cons. are selected based on their representativeness. Detailed discussion can be found in Appendix A.
@@ 2.1 Large Language models (LLMs)
Large language models (LLMs) pre-trained on large-scale corpus have shown great potential in various NLP tasks [13]. As shown in Fig. 3, most LLMs derive from the Transformer design [50], which contains the encoder and decoder modules empowered by a self-attention mechanism. Based on the architecture structure, LLMs can be categorized into three groups: 1) encoder-only LLMs2) encoder-decoder LLMs, and 3) decoder-only LLMs. As shown in Fig. 2, we summarize several representative LLMs with different model architectures, model sizes, and open-source availabilities.
@@ 2.1.4 Prompt Engineering
Prompt engineering seeks to improve the capacity of large large language models (e.g., ChatGPT) in diverse complex tasks such as question answering, sentiment classification, and common sense reasoning. Chain-of-thought (CoT) prompt [63] enables complex reasoning capabilities through intermediate reasoning steps. Prompt engineering also enables the integration of structural data like knowledge graphs (KGs) into LLMs. Li et al. [64] simply linearizes the KGs and uses templates to convert the KGs into passages. Mindmap [65] designs a KG prompt to convert graph structure into a mind map that enables LLMs to perform reasoning on it. Prompt offers a simple way to utilize the potential of LLMs without finetuning. Proficiency in prompt engineering leads to a better understanding of the strengths and weaknesses of LLMs.
@@ 2.2.3 Domain-specific Knowledge Graphs
OpenBG Recommendation https://shorturl.at/pDMV9
@@ 2.3 Applications
LLMs as KGs have been widely applied in various real-world applications. We summarize some representative applications of using LLMs and KGs in Table I. ChatGPT/GPT-4 are LLM-based chatbots that can communicate with humans in a natural dialogue format. To improve knowledge awareness of LLMs, ERNIE 3.0 and Bard incorporate KGs into their chatbot applications. Instead of Chatbot. Firefly develops a photo editing application that allows users to edit photos by using natural language descriptions. Copilot, New Bing, and Shop.ai adopt LLMs to empower their applications in the areas of coding assistant, web search, and recommendation, respectively. Wikidata and KO are two representative knowledge graph applications that are used to provide external knowledge. OpenBG [90] is a knowledge graph designed for recommendation. Doctor.ai develops a health care assistant that incorporates LLMs and KGs to provide medical advice.
@@ 3.1 Roadmap
The roadmap of unifying KGs and LLMs is illustrated in Fig. 6. In the roadmap, we identify three frameworks for the unification of LLMs and KGs, including KG-enhanced LLMs, LLM-augmented KGs, and Synergized LLMs + KGs. The KG-enhanced LLMs and LLM-augmented KGs are two parallel frameworks that aim to enhance the capabilities of LLMs and KGs, respectively. Building upon these frameworks, Synergized LLMs + KGs is a unified framework that aims to synergize LLMs and KGs to mutually enhance each other.
@@ 4 KG-enhanced LLMs
[b] Task Method Year KG Technique KG-enhanced LLM pre-training ERNIE [35] 2019 E Integrating KGs into Training Objective GLM [102] 2020 C Integrating KGs into Training Objective Ebert [103] 2020 D Integrating KGs into Training Objective KEPLER [40] 2021 E Integrating KGs into Training Objective Deterministic LLM [104] 2022 E Integrating KGs into Training Objective KALA [105] 2022 D Integrating KGs into Training Objective WKLM [106] 2020 E Integrating KGs into Training Objective K-BERT [36] 2020 E + D Integrating KGs into Language Model Inputs CoLAKE [107] 2020 E Integrating KGs into Language Model Inputs ERNIE3.0 [101] 2021 E + D Integrating KGs into Language Model Inputs DkLLM [108] 2022 E Integrating KGs into Language Model Inputs KP-PLM [109] 2022 E KGs Instruction-tuning OntoPrompt [110] 2022 E + D KGs Instruction-tuning ChatKBQA [111] 2023 E KGs Instruction-tuning RoG [112] 2023 E KGs Instruction-tuning KG-enhanced LLM inference KGLM [113] 2019 E Retrival-augmented knowledge fusion REALM [114] 2020 E Retrival-augmented knowledge fusion RAG [92] 2020 E Retrival-augmented knowledge fusion EMAT [115] 2022 E Retrival-augmented knowledge fusion Li et al. [64] 2023 C KGs Prompting Mindmap [65] 2023 E + D KGs Prompting ChatRule [116] 2023 E + D KGs Prompting CoK [117] 2023 E + C + D KGs Prompting KG-enhanced LLM interpretability LAMA [14] 2019 E KGs for LLM probing LPAQA [118] 2020 E KGs for LLM probing Autoprompt [119] 2020 E KGs for LLM probing MedLAMA [120] 2022 D KGs for LLM probing LLM-facteval [121] 2023 E + D KGs for LLM probing KagNet [38] 2019 C KGs for LLM analysis Interpret-lm [122] 2021 E KGs for LLM analysis knowledge-neurons [39] 2021 E KGs for LLM analysis Shaobo et al. [123] 2022 E KGs for LLM analysis
  • E: Encyclopedic Knowledge Graphs, C: Commonsense Knowledge Graphs, D: Domain-Specific Knowledge Graphs.
@@ 4.1 KG-enhanced LLM Pre-training
Existing large language models mostly rely on unsupervised training on the large-scale corpus. While these models may exhibit impressive performance on downstream tasks, they often lack practical knowledge relevant to the real world. Previous works that integrate KGs into large language models can be categorized into three parts: 1) Integrating KGs into training objective, 2) Integrating KGs into LLM inputs, and 3) KGs Instruction-tuning.
@@ 4.1.1 Integrating KGs into Training Objective
The other line of work explicitly leverages the connections with knowledge and input text. As shown in Fig. 9, ERNIE [35] proposes a novel word-entity alignment training objective as a pre-training objective. Specifically, ERNIE feeds both sentences and corresponding entities mentioned in the text into LLMs, and then trains the LLMs to predict alignment links between textual tokens and entities in knowledge graphs. Similarly, KALM [91] enhances the input tokens by incorporating entity embeddings and includes an entity prediction pre-training task in addition to the token-only pre-training objective. This approach aims to improve the ability of LLMs to capture knowledge related to entities. Finally, KEPLER [40] directly employs both knowledge graph embedding training objective and Masked token pre-training objective into a shared transformer-based encoder. Deterministic LLM [104] focuses on pre-training language models to capture deterministic factual knowledge. It only masks the span that has a deterministic entity as the question and introduces additional clue contrast learning and clue classification objective. WKLM [106] first replaces entities in the text with other same-type entities and then feeds them into LLMs. The model is further pre-trained to distinguish whether the entities have been replaced or not.
@@ 4.1.2 Integrating KGs into LLM Inputs
The above methods can indeed inject a large amount of knowledge into LLMs. However, they mostly focus on popular entities and overlook the low-frequent and long-tail ones. DkLLM [108] aims to improve the LLMs representations towards those entities. DkLLM first proposes a novel measurement to determine long-tail entities and then replaces these selected entities in the text with pseudo token embedding as new input to the large language models. Furthermore, Dict-BERT [125] proposes to leverage external dictionaries to solve this issue. Specifically, Dict-BERT improves the representation quality of rare words by appending their definitions from the dictionary at the end of input text and trains the language model to locally align rare word representations in input sentences and dictionary definitions as well as to discriminate whether the input text and definition are correctly mapped.
@@
4.1.3 KGs Instruction-tuning
Instead of injecting factual knowledge into LLMs, the KGs Instruction-tuning aims to fine-tune LLMs to better comprehend the structure of KGs and effectively follow user instructions to conduct complex tasks. KGs Instruction-tuning utilizes both facts and the structure of KGs to create instruction-tuning datasets. LLMs finetuned on these datasets can extract both factual and structural knowledge from KGs, enhancing the reasoning ability of LLMs. KP-PLM [109] first designs several prompt templates to transfer structural graphs into natural language text. Then, two self-supervised tasks are proposed to finetune LLMs to further leverage the knowledge from these prompts. OntoPrompt [110] proposes an ontology-enhanced prompt-tuning that can place knowledge of entities into the context of LLMs, which are further finetuned on several downstream tasks. ChatKBQA [111] finetunes LLMs on KG structure to generate logical queries, which can be executed on KGs to obtain answers. To better reason on graphs, RoG [112] presents a planning-retrieval-reasoning framework. RoG is finetuned on KG structure to generate relation paths grounded by KGs as faithful plans. These plans are then used to retrieve valid reasoning paths from the KGs for LLMs to conduct faithful reasoning and generate interpretable results.
KGs Instruction-tuning can better leverage the knowledge from KGs for downstream tasks. However, it requires retraining the models, which is time-consuming and requires lots of resources.
@@ 4.2 KG-enhanced LLM Inference
The above methods could effectively fuse knowledge into LLMs. However, real-world knowledge is subject to change and the limitation of these approaches is that they do not permit updates to the incorporated knowledge without retraining the model. As a result, they may not generalize well to the unseen knowledge during inference [126]. Therefore, considerable research has been devoted to keeping the knowledge space and text space separate and injecting the knowledge while inference. These methods mostly focus on the Question Answering (QA) tasks, because QA requires the model to capture both textual semantic meanings and up-to-date real-world knowledge.
@@ 4.2.1 Retrieval-Augmented Knowledge Fusion
Retrieval-Augmented Knowledge Fusion is a popular method to inject knowledge into LLMs during inference. The key idea is to retrieve relevant knowledge from a large corpus and then fuse the retrieved knowledge into LLMs. As shown in Fig. 11, RAG [92] proposes to combine non-parametric and parametric modules to handle the external knowledge. Given the input text, RAG first searches for relevant KG in the non-parametric module via MIPS to obtain several documents. RAG then treats these documents as hidden variables $z$ and feeds them into the output generator, empowered by Seq2Seq LLMs, as additional context information. The research indicates that using different retrieved documents as conditions at different generation steps performs better than only using a single document to guide the whole generation process. The experimental results show that RAG outperforms other parametric-only and non-parametric-only baseline models in open-domain QA. RAG can also generate more specific, diverse, and factual text than other parameter-only baselines. Story-fragments [127] further improves architecture by adding an additional module to determine salient knowledge entities and fuse them into the generator to improve the quality of generated long stories. EMAT [115] further improves the efficiency of such a system by encoding external knowledge into a key-value memory and exploiting the fast maximum inner product search for memory querying. REALM [114] proposes a novel knowledge retriever to help the model to retrieve and attend over documents from a large corpus during the pre-training stage and successfully improves the performance of open-domain question answering. KGLM [113] selects the facts from a knowledge graph using the current context to generate factual sentences. With the help of an external knowledge graph, KGLM could describe facts using out-of-domain words or phrases.
@@
4.2.2 KGs Prompting
To better feed the KG structure into the LLM during inference, KGs prompting aims to design a crafted prompt that converts structured KGs into text sequences, which can be fed as context into LLMs. In this way, LLMs can better take advantage of the structure of KGs to perform reasoning. Li et al. [64] adopt the pre-defined template to convert each triple into a short sentence, which can be understood by LLMs for reasoning. Mindmap [65] designs a KG prompt to convert graph structure into a mind map that enables LLMs to perform reasoning by consolidating the facts in KGs and the implicit knowledge from LLMs. ChatRule [116] samples several relation paths from KGs, which are verbalized and fed into LLMs. Then, LLMs are prompted to generate meaningful logical rules that can be used for reasoning. CoK [117] proposes a chain-of-knowledge prompting that uses a sequence of triples to elicit the reasoning ability of LLMs to reach the final answer.
KGs prompting presents a simple way to synergize LLMs and KGs. By using the prompt, we can easily harness the power of LLMs to perform reasoning based on KGs without retraining the models. However, the prompt is usually designed manually, which requires lots of human effort.
4.3 Comparison between KG-enhanced LLM Pre-training and Inference
KG-enhanced LLM Pre-training methods commonly enrich large-amount of unlabeled corpus with semantically relevant real-world knowledge. These methods allow the knowledge representations to be aligned with appropriate linguistic context and explicitly train LLMs to leverage those knowledge from scratch. When applying the resulting LLMs to downstream knowledge-intensive tasks, they should achieve optimal performance. In contrast, KG-enhanced LLM inference methods only present the knowledge to LLMs in the inference stage and the underlying LLMs may not be trained to fully leverage these knowledge when conducting downstream tasks, potentially resulting in sub-optimal model performance.
However, real-world knowledge is dynamic and requires frequent updates. Despite being effective, the KG-enhanced LLM Pre-training methods never permit knowledge updates or editing without model re-training. As a result, the KG-enhanced LLM Pre-training methods could generalize poorly to recent or unseen knowledge. KG-enhanced LLM inference methods can easily maintain knowledge updates by changing the inference inputs. These methods help improve LLMs performance on new knowledge and domains.
In summary, when to use these methods depends on the application scenarios. If one wishes to apply LLMs to handle time-insensitive knowledge in particular domains (e.g., commonsense and reasoning knowledge), KG-enhanced LLM Pre-training methods should be considered. Otherwise, KG-enhanced LLM inference methods can be used to handle open-domain knowledge with frequent updates.